# Deep Learning Meets Mechanism Design: A Survey

### v2

@@ IV-A RochetNet
An optimal auction design that maximizes the expected revenue to the seller under IC is proposed in [DUTTING23]. This approach is a characterization based approach where IC constraints are encoded within the Neural Network (NN) architecture. The authors call such an NN a RochetNet. This setup is for a single buyer with additive valuations and a seller with multiple items to sell. Let $\textbf{u}^{\textbf{w},\textbf{w}_{0}}(\textbf{v})$ be the seller’s utility function, defined as
@@
Figure 2: RochetNet: NN implementation of Eq. 16 (Figure reproduced from [DUTTING23])
Figure 3: Example of utilization function generated from RochetNet (Figure reproduced from [DUTTING23])
@@ IV-B RegretNet
The next model in [DUTTING23] uses differentiable approximation for IC constraints and the objective function is augmented with the term that measures how much IC constraints are violated. This approach is called a characterization free approach and is applicable to a mechanism where a seller with multiple items wishes to sell to multiple buyers. The authors define ex post regret for each buyer $i$ as:
@@
Figure 4: RegretNet: The allocation and payment networks for $n$ additive buyers with $m$ items (Figure reproduced from [DUTTING23])
@@ IV-C MyersonNet
[DUTTING23] also proposes an NN architecture for a seller with a single item (by letting $m=1$ in their proposed setting) and multiple buyers and extends the work to a general setting with multiple buyers and multiple items with combinatorial valuations. Below, we provide the NN architecture for a single item with multiple buyers. In this setup, the valuation function becomes $\textbf{v}=[v_{1},\dots v_{n}]$ for $n$ buyers with $\textbf{b}\leq\textbf{v}$.Then, the utility of the $i^{th}$ buyer is
@@
Consider a distribution function $F_{i}$ having probability density $f_{i}$, then the virtual valuation function is given by $\xi_{i}(v_{i})=v_{i}-\frac{1-F(v_{i})}{f(v_{i})}$ and for monotonically non-decreasing $\xi_{i}$, the corresponding distribution $F_{i}$ (and density $f_{i}$) is regular. This implies that the virtual valuation, $\bar{\Phi}_{i}=\xi_{i}$ for regular distributions $F_{i}\;\forall i=1,2,\dots,n$. If these virtual functions $\xi_{i}$ are monotonically increasing, then the optimal auction can be designed by first applying monotone transformations to bids and then feeding these virtual values to the second price auction with zero reserve price, leading to an output $(P^{o},t^{o})$ with allocation $P^{o}(\bar{b})$ and payment $\bar{\Phi}_{i}^{-1}(t^{o}_{i}(\bar{b}))$. This auction mechanism is DSIC [DUTTING23], [ROCHET87]. These virtual value functions $\bar{\Phi}_{i}$ are modelled using two layer feed forward network with min and max over linear functions for $K$ groups of $J$ linear functions as shown below.
@@
Figure 5: Monotone virtual value function implementation given in Eq. 22 where $h_{k,j}(b_{i})=e^{\alpha^{i}_{k,j}}b_{i}+\beta^{i}_{k,j}$ (Figure reproduced from [DUTTING23])
Figure 6: MyersonNet: monotone transformed bids are applied to Second Price auction network (Figure reproduced from [DUTTING23])
@@
Figure 7: Second Price auction with zero reserve price network (Figure reproduced from [DUTTING23])
@@ IV-D MenuNet
In [SHEN18], the authors propose an NN based optimal auction design by introducing another independent NN structure that captures the behaviour of buyers. This buyer network models the exact IC constraint. The authors propose an NN model for an auction setting where a seller with multiple heterogeneous ($m$ items) items would like to sell to a single buyer who may have an additive valuation function.
Let p be the allocation vector and $\textbf{p}\in\mathscr{P}\subseteq[0,1]^{m}$ where $p_{i}$ is the probability of allocating an item $i$ to the buyer. Let $\textbf{v}=(v_{1},\dots,v_{m})$ be a value profile where $v_{i}\sim F_{i}$ is the value of an $i^{th}$ item with $0\leq v_{i}\leq\hat{v}_{i}$ and let $V_{i}$ be the set of $d_{i}$ discrete values of $v_{i}$ in the interval $0\leq v_{i}\leq\hat{v}_{i}$. Let us define $V=V_{1}\times,\dots,\times V_{m}$, $\mathscr{F}=F_{1}\times,\dots,\times F_{m}$ to be independent in each dimension and $t$ to be the payment. The proposed MenuNet architecture is shown in Fig. 8. This architecture is based on the Naive mechanism ([SHEN18]) which shows allocations along with corresponding items known as the menu and the buyer makes the most preferred choice. The high level structure of the Naive mechanism and the Direct mechanism is shown in Figs. 9, 10. As shown in the figures, in the case of direct mechanisms, the objective is evaluated on the outcomes that are restricted by IC and IR constraints whereas in the case of Naive mechanisms, the mechanism’s outcome serves as input to the buyer choices, which are used to compute the objective.
@@
Figure 8: MENUNet Architecture (Figure reproduced from [SHEN18])
@@
Figure 9: Naive Mechanism Architecture (Figure reproduced from [SHEN18])
Figure 10: Direct Mechanism Architecture (Figure reproduced from [SHEN18])
@@ IV-E RegretFormer
In [IVANOV22], the authors provide non-trivial extensions to the RegretNet architecture [DUTTING23] in two ways. The first is, they propose a novel NN architecture based on attention layers which they call RegretFormer, shown in Fig. 11. The RegretNet architecture has three main issues. It is not generalizeable to unseen number of items and buyers or to data with heterogeneous input sizes, it is sensitive to the order of items and buyers in the input bid matrix, and the fully connected layers alone do not produce the expressive power needed for auction design. RegretFormer tackles the latter 2 problems.
Figure 11: RegretFormer Architecture in (Figure reproduced from [IVANOV22])
The second way they improve upon RegretNet is by attempting to maximize the revenue subject to a maximum regret cap. RegretNet aims to simultaneously maximize revenue and minimize regret by finding suitable values for the hyperparameters of the model. This technique imposes extremely precise hyperparameters for every experiment which have no predictability, i.e., selecting a set of values for the hyperparameters will not guarantee any outcome. This also means that whenever a new experiment must be conducted, for instance with a different number of items and buyers, hyperparameter tuning will need to be extensively performed to obtain accurate results. [IVANOV22] maximizes the revenue given a maximum regret budget $R_{max}$ which is specified by the designer beforehand. This reformulates the problem in Eq. 19 as
@@
The dual gradient descent technique [BOYD04] is applied by performing one gradient update for the parameters w and one gradient update for $\gamma$.
@@ IV-F Budgeted RegretNet
There are scenarios where buyers have budget constraints in purchasing an item such as spectrum auctions, land auctions, etc. This constraint specifies that even though buyers know the exact value of the item, they are not able to pay that much money due to their financial constraints. This type of auction has the valuation of an item along with the budget constraint as private information of buyers. In [FENG18], the authors propose a DL based approach for maximizing revenue under the private budget constraints that satisfy both BIC and conditional IC constraints by extending the RegretNet framework [DUTTING23] which they call Budgeted RegretNet. Let $T_{i}=(B_{i},v_{i})\sim F_{i}$ be the possible type space for buyer $i$, where $B_{i}$, $v_{i}$, are the payment budget and valuation for the $i^{th}$ buyer of a single item respectively (drawn from distribution $F_{i}$ and $\textbf{T}=(T_{1},\dots,T_{n})\sim F=\prod_{i=1}^{n}F_{i}$), and sold by a single seller where this information is private to buyer $i$. Every buyer bids with $t^{\prime}\in T_{i}$ and the utility of each buyer $i$ given in equation 20, changes to
@@ IV-G Stage-IC and Dynamic-IC
[DENG20] proposes a new metric for quantifying IC, using a data driven computational approach. It is a measure of the marginal advantage of bidding a scaled version of the true value. This metric can be applied to both static and dynamic auctions. In dynamic auctions, agents’ previous bids influence their future payments. The metric is described below.
@@ V Learning Welfare Maximizing Mechanisms
[TACCHETTI19] focuses on DL based approaches for auctions that incentivize truthful bids, maximize social welfare, and minimize the economic burden on buyers. Theoretically, all these properties cannot be simultaneously satisfied. Hence, the authors design Neural Networks (NNs) to approximately achieve these properties. The NN architecture consists of Convolutional Neural Networks (CNNs) that learn Groves’ payment rule for truthful mechanism. The NN architecture is proposed for auctions with multiple units with decreasing marginal utilities. [BRERO19] proposes a Machine Learning (ML) based Combinatorial Auction (CA) that improves preference elicitation based on value queries. In this method, each buyer’s value function is learned via separate ML models (bundle-value pair as input) and these valuations are fed to another ML model that has the winner determination algorithm. The overall goal of the proposed model is to maximize the learned social welfare of all buyers. In [WEISSTEINER20] the authors propose DL based iterative CAs where they first formulate a Deep Neural Network (DNN) based Winner Determination Problem as a mixed integer program problem and then compare the results with Support Vector Regression (SVR).
The authors in [DENG22] propose obtaining reserve prices through ML techniques, which can be given as advice to participating buyers. They show that through this advice, buyers with large total values across auctions are limited in their power to manipulate the outcome of the auction. They also provide individual welfare bound on VCG auctions where reserve prices are approximated through ML techniques.
In the case of spectrum auctions, the main goal of the auctioneer is to maximize the total welfare of the participants, to create more jobs, to maximize connectivity, and to increase trade. [TACCHETTI19] studies an efficient and truthful multi-unit auction Mechanism Design using NNs. It models multi-unit auctions as a VCG auction mechanism and proposes a model that uses the CNN architecture with loss function as the payment calculated using Groves’ payment rule. The main objective of the paper is to incentivize truthful buyers and allocate goods that maximize the social welfare of all the buyers and minimize the economic burden on the buyers. This is achieved via designing truthful and efficient auctions that minimizes the sum of payments under the constraints of IR and weak budget balance and choosing the Groves payment rule by selecting a function $h:\textbf{V}_{-i}\rightarrow R$. The mathematical formulation is given below. For a set $M$ of $m$ items and a set $N$ of $n$ buyers, let $K$ be the set that consists of all possible ways to allocate the items to buyers. The utility of the buyer $i$ is $U_{i}(k,t_{i})=v_{i}(k)-t_{i}$ where $v_{i}(k)$ is the valuation of the $k^{th}$ allocation by user $i$ and $t_{i}$ is the payment for user $i$.
@@
Figure 12: Multi-item auction with social welfare optimization using CNN (Figure reproduced from [TACCHETTI19])
@@
CAs such as spectrum auctions and allocation of ad-slots in TV are used to allocate multiple items in a bundle to multiple buyers. The main challenge is that when the number of items grow, the bundle size also grows exponentially. This causes a more challenging preference elicitation. [BRERO19] proposes an auction mechanism that uses a traditional ML technique, known as Machine Learning Powered Iterative Combinatorial Auction (MLCA) with the goal of maximizing empirical efficiency in a realistic CA. In this architecture, an ML algorithm for each buyer is trained on value reports generalizable to the whole bundle space. In every round of an auction, the MLCA algorithm uses ML-powered query sub-routines that contain the learned valuations for each buyer and the winner determination steps. The mathematical formulation of the model is given below.
@@
Figure 13: Schematic Representation of Query module (Figure reproduced from [BRERO19])
@@
The MLCA algorithm uses the query module sub-routine to compute feasible allocations and obtains the bids from the buyers through an iterative process. From the collected bids, it computes the final allocation and each buyers’ payments using the equations given above in the outcome computation rules. As opposed to SVR for the winner determination problem in [BRERO19], the authors of [WEISSTEINER20] use a DNN based winner determination problem for the Iterative CA that has the same problem formulation as that of [BRERO19]. Let $\hat{v}^{t}_{i}$ be the true value estimator of an $i^{th}$ buyer and the data set to be trained be $B^{t-1}_{i}$ (i.e,. queried values are up to round t1)t-1). Then, the estimated social welfare function is
@@
Figure 14: Representation of a DNN $\textit{N}_{i}$ (Figure reproduced from [WEISSTEINER20])
@@ VI Learning Mechanisms with Fairness in Allocation
[KUO20] introduces a fairness condition in addition to revenue maximization and incentive compatibility to design an auction using a DL model (ProportionNet). This technique is an extension to RegretNet proposed in [DUTTING23]. [BARMAN19] studies Envy freeness up to one good (EF1), maximin share guarantee (MMS) and Nash social welfare (NSW) fairness conditions. Using these metrics, [AHMAD21] proposes an NN model for an auction whose allocation is fair as well as IC and IR while maximizing the social welfare metric.
There are some scenarios wherein the auctioneer may be interested in the fair allocation of resources to buyers, for example, the allocation of resources such as spectrum allocation and the online advertisement of career opportunities to all the users irrespective of categories. [KUO20] studies revenue maximizing auctions that have strategy proof and fairness constraints for online advertisement auction scenarios. It extends the work of RegretNet studied in [DUTTING23]. First, let us define one of the fairness metrics known as total variation of fairness [ILVENTO20].
@@
The fairness constraint is ensured by adding the loss function $L_{unf}$ with the loss function defined in Eq. 36 for the RegretNet model (Fig. 4) [DUTTING23] using the augmented Lagrangian approach.
In many real world applications we consider the fair allocation of resources, for example, in course assignments, ad-auctions, and inventory pricing. In [BARMAN19], the authors consider the auctioning of indivisible goods to multiple strategic buyers in a single-parameter environment. They study EF1, MMS, and NSW fairness conditions. Let there be $n$ buyers and a single seller having $m$ goods to sell. Let $v_{i}\in R_{+}$ be the valuation for buyer $i$ (private value) and let the known public value summarization function among all buyers be $W:2^{[m]}\rightarrow R_{+}$. For a given subset $S\subseteq[m]$, the valuation of the $i^{th}$ buyer is $v_{i}(S)=v_{i}W(S)$. Then, the fairness conditions are defined below.
@@
Motivated by the fairness metrics of [BARMAN19], [AHMAD21] proposes an NN model for an auction that has a fair allocation in addition to being IC and IR while maximizing social welfare. The proposed model is described below.

### v3

@@ IV-A RochetNet
An optimal auction design that maximizes the expected revenue to the seller under IC is proposed in [14]. This approach is a characterization based approach where IC constraints are encoded within the Neural Network (NN) architecture. The authors call such an NN a RochetNet. This setup is for a single buyer with additive valuations and a seller with multiple items to sell. Let $\textbf{u}^{\textbf{w},\textbf{w}_{0}}(\textbf{v})$ be the seller’s utility function, defined as
@@
Figure 2: RochetNet: NN implementation of Eq. 16 (Figure reproduced from [14])
Figure 3: Example of utilization function generated from RochetNet (Figure reproduced from [14])
@@ IV-B RegretNet
The next model in [14] uses differentiable approximation for IC constraints and the objective function is augmented with the term that measures how much IC constraints are violated. This approach is called a characterization free approach and is applicable to a mechanism where a seller with multiple items wishes to sell to multiple buyers. The authors define ex post regret for each buyer $i$ as:
@@
Figure 4: RegretNet: The allocation and payment networks for $n$ additive buyers with $m$ items (Figure reproduced from [14])
@@ IV-C MyersonNet
[14] also proposes an NN architecture for a seller with a single item (by letting $m=1$ in their proposed setting) and multiple buyers and extends the work to a general setting with multiple buyers and multiple items with combinatorial valuations. Below, we provide the NN architecture for a single item with multiple buyers. In this setup, the valuation function becomes $\textbf{v}=[v_{1},\dots v_{n}]$ for $n$ buyers with $\textbf{b}\leq\textbf{v}$.Then, the utility of the $i^{th}$ buyer is
@@
Consider a distribution function $F_{i}$ having probability density $f_{i}$, then the virtual valuation function is given by $\xi_{i}(v_{i})=v_{i}-\frac{1-F(v_{i})}{f(v_{i})}$ and for monotonically non-decreasing $\xi_{i}$, the corresponding distribution $F_{i}$ (and density $f_{i}$) is regular. This implies that the virtual valuation, $\bar{\Phi}_{i}=\xi_{i}$ for regular distributions $F_{i}\;\forall i=1,2,\dots,n$. If these virtual functions $\xi_{i}$ are monotonically increasing, then the optimal auction can be designed by first applying monotone transformations to bids and then feeding these virtual values to the second price auction with zero reserve price, leading to an output $(P^{o},t^{o})$ with allocation $P^{o}(\bar{b})$ and payment $\bar{\Phi}_{i}^{-1}(t^{o}_{i}(\bar{b}))$. This auction mechanism is DSIC [14], [42]. These virtual value functions $\bar{\Phi}_{i}$ are modelled using two layer feed forward network with min and max over linear functions for $K$ groups of $J$ linear functions as shown below.
@@
Figure 5: Monotone virtual value function implementation given in Eq. 22 where $h_{k,j}(b_{i})=e^{\alpha^{i}_{k,j}}b_{i}+\beta^{i}_{k,j}$ (Figure reproduced from [14])
Figure 6: MyersonNet: monotone transformed bids are applied to Second Price auction network (Figure reproduced from [14])
@@
Figure 7: Second Price auction with zero reserve price network (Figure reproduced from [14])
@@ IV-D MenuNet
In [43], the authors propose an NN based optimal auction design by introducing another independent NN structure that captures the behaviour of buyers. This buyer network models the exact IC constraint. The authors propose an NN model for an auction setting where a seller with multiple heterogeneous ($m$ items) items would like to sell to a single buyer who may have an additive valuation function.
Let p be the allocation vector and $\textbf{p}\in\mathscr{P}\subseteq[0,1]^{m}$ where $p_{i}$ is the probability of allocating an item $i$ to the buyer. Let $\textbf{v}=(v_{1},\dots,v_{m})$ be a value profile where $v_{i}\sim F_{i}$ is the value of an $i^{th}$ item with $0\leq v_{i}\leq\hat{v}_{i}$ and let $V_{i}$ be the set of $d_{i}$ discrete values of $v_{i}$ in the interval $0\leq v_{i}\leq\hat{v}_{i}$. Let us define $V=V_{1}\times,\dots,\times V_{m}$, $\mathscr{F}=F_{1}\times,\dots,\times F_{m}$ to be independent in each dimension and $t$ to be the payment. The proposed MenuNet architecture is shown in Fig. 8. This architecture is based on the Naive mechanism ([43]) which shows allocations along with corresponding items known as the menu and the buyer makes the most preferred choice. The high level structure of the Naive mechanism and the Direct mechanism is shown in Figs. 9, 10. As shown in the figures, in the case of direct mechanisms, the objective is evaluated on the outcomes that are restricted by IC and IR constraints whereas in the case of Naive mechanisms, the mechanism’s outcome serves as input to the buyer choices, which are used to compute the objective.
@@
Figure 8: MENUNet Architecture (Figure reproduced from [43])
@@
Figure 9: Naive Mechanism Architecture (Figure reproduced from [43])
Figure 10: Direct Mechanism Architecture (Figure reproduced from [43])
@@ IV-E RegretFormer
In [21], the authors provide non-trivial extensions to the RegretNet architecture [14] in two ways. The first is, they propose a novel NN architecture based on attention layers which they call RegretFormer, shown in Fig. 11. The RegretNet architecture has three main issues. It is not generalizeable to unseen number of items and buyers or to data with heterogeneous input sizes, it is sensitive to the order of items and buyers in the input bid matrix, and the fully connected layers alone do not produce the expressive power needed for auction design. RegretFormer tackles the latter 2 problems.
Figure 11: RegretFormer Architecture in (Figure reproduced from [21])
The second way they improve upon RegretNet is by attempting to maximize the revenue subject to a maximum regret cap. RegretNet aims to simultaneously maximize revenue and minimize regret by finding suitable values for the hyperparameters of the model. This technique imposes extremely precise hyperparameters for every experiment which have no predictability, i.e., selecting a set of values for the hyperparameters will not guarantee any outcome. This also means that whenever a new experiment must be conducted, for instance with a different number of items and buyers, hyperparameter tuning will need to be extensively performed to obtain accurate results. [21] maximizes the revenue given a maximum regret budget $R_{max}$ which is specified by the designer beforehand. This reformulates the problem in Eq. 19 as
@@
The dual gradient descent technique [6] is applied by performing one gradient update for the parameters w and one gradient update for $\gamma$.
@@ IV-F Budgeted RegretNet
There are scenarios where buyers have budget constraints in purchasing an item such as spectrum auctions, land auctions, etc. This constraint specifies that even though buyers know the exact value of the item, they are not able to pay that much money due to their financial constraints. This type of auction has the valuation of an item along with the budget constraint as private information of buyers. In [15], the authors propose a DL based approach for maximizing revenue under the private budget constraints that satisfy both BIC and conditional IC constraints by extending the RegretNet framework [14] which they call Budgeted RegretNet. Let $T_{i}=(B_{i},v_{i})\sim F_{i}$ be the possible type space for buyer $i$, where $B_{i}$, $v_{i}$, are the payment budget and valuation for the $i^{th}$ buyer of a single item respectively (drawn from distribution $F_{i}$ and $\textbf{T}=(T_{1},\dots,T_{n})\sim F=\prod_{i=1}^{n}F_{i}$), and sold by a single seller where this information is private to buyer $i$. Every buyer bids with $t^{\prime}\in T_{i}$ and the utility of each buyer $i$ given in equation 20, changes to
@@ IV-G Stage-IC and Dynamic-IC
[12] proposes a new metric for quantifying IC, using a data driven computational approach. It is a measure of the marginal advantage of bidding a scaled version of the true value. This metric can be applied to both static and dynamic auctions. In dynamic auctions, agents’ previous bids influence their future payments. The metric is described below.
@@ V Learning Welfare Maximizing Mechanisms
[44] focuses on DL based approaches for auctions that incentivize truthful bids, maximize social welfare, and minimize the economic burden on buyers. Theoretically, all these properties cannot be simultaneously satisfied. Hence, the authors design Neural Networks (NNs) to approximately achieve these properties. The NN architecture consists of Convolutional Neural Networks (CNNs) that learn Groves’ payment rule for truthful mechanism. The NN architecture is proposed for auctions with multiple units with decreasing marginal utilities. [7] proposes a Machine Learning (ML) based Combinatorial Auction (CA) that improves preference elicitation based on value queries. In this method, each buyer’s value function is learned via separate ML models (bundle-value pair as input) and these valuations are fed to another ML model that has the winner determination algorithm. The overall goal of the proposed model is to maximize the learned social welfare of all buyers. In [46] the authors propose DL based iterative CAs where they first formulate a Deep Neural Network (DNN) based Winner Determination Problem as a mixed integer program problem and then compare the results with Support Vector Regression (SVR).
The authors in [11] propose obtaining reserve prices through ML techniques, which can be given as advice to participating buyers. They show that through this advice, buyers with large total values across auctions are limited in their power to manipulate the outcome of the auction. They also provide individual welfare bound on VCG auctions where reserve prices are approximated through ML techniques.
In the case of spectrum auctions, the main goal of the auctioneer is to maximize the total welfare of the participants, to create more jobs, to maximize connectivity, and to increase trade. [44] studies an efficient and truthful multi-unit auction Mechanism Design using NNs. It models multi-unit auctions as a VCG auction mechanism and proposes a model that uses the CNN architecture with loss function as the payment calculated using Groves’ payment rule. The main objective of the paper is to incentivize truthful buyers and allocate goods that maximize the social welfare of all the buyers and minimize the economic burden on the buyers. This is achieved via designing truthful and efficient auctions that minimizes the sum of payments under the constraints of IR and weak budget balance and choosing the Groves payment rule by selecting a function $h:\textbf{V}_{-i}\rightarrow R$. The mathematical formulation is given below. For a set $M$ of $m$ items and a set $N$ of $n$ buyers, let $K$ be the set that consists of all possible ways to allocate the items to buyers. The utility of the buyer $i$ is $U_{i}(k,t_{i})=v_{i}(k)-t_{i}$ where $v_{i}(k)$ is the valuation of the $k^{th}$ allocation by user $i$ and $t_{i}$ is the payment for user $i$.
@@
Figure 12: Multi-item auction with social welfare optimization using CNN (Figure reproduced from [44])
@@
CAs such as spectrum auctions and allocation of ad-slots in TV are used to allocate multiple items in a bundle to multiple buyers. The main challenge is that when the number of items grow, the bundle size also grows exponentially. This causes a more challenging preference elicitation. [7] proposes an auction mechanism that uses a traditional ML technique, known as Machine Learning Powered Iterative Combinatorial Auction (MLCA) with the goal of maximizing empirical efficiency in a realistic CA. In this architecture, an ML algorithm for each buyer is trained on value reports generalizable to the whole bundle space. In every round of an auction, the MLCA algorithm uses ML-powered query sub-routines that contain the learned valuations for each buyer and the winner determination steps. The mathematical formulation of the model is given below.
@@
Figure 13: Schematic Representation of Query module (Figure reproduced from [7])
@@
The MLCA algorithm uses the query module sub-routine to compute feasible allocations and obtains the bids from the buyers through an iterative process. From the collected bids, it computes the final allocation and each buyers’ payments using the equations given above in the outcome computation rules. As opposed to SVR for the winner determination problem in [7], the authors of [46] use a DNN based winner determination problem for the Iterative CA that has the same problem formulation as that of [7]. Let $\hat{v}^{t}_{i}$ be the true value estimator of an $i^{th}$ buyer and the data set to be trained be $B^{t-1}_{i}$ (i.e,. queried values are up to round t1)t-1). Then, the estimated social welfare function is
@@
Figure 14: Representation of a DNN $\textit{N}_{i}$ (Figure reproduced from [46])
@@ VI Learning Mechanisms with Fairness in Allocation
[24] introduces a fairness condition in addition to revenue maximization and incentive compatibility to design an auction using a DL model (ProportionNet). This technique is an extension to RegretNet proposed in [14]. [3] studies Envy freeness up to one good (EF1), maximin share guarantee (MMS) and Nash social welfare (NSW) fairness conditions. Using these metrics, [1] proposes an NN model for an auction whose allocation is fair as well as IC and IR while maximizing the social welfare metric.
There are some scenarios wherein the auctioneer may be interested in the fair allocation of resources to buyers, for example, the allocation of resources such as spectrum allocation and the online advertisement of career opportunities to all the users irrespective of categories. [24] studies revenue maximizing auctions that have strategy proof and fairness constraints for online advertisement auction scenarios. It extends the work of RegretNet studied in [14]. First, let us define one of the fairness metrics known as total variation of fairness [20].
@@
The fairness constraint is ensured by adding the loss function $L_{unf}$ with the loss function defined in Eq. 36 for the RegretNet model (Fig. 4) [14] using the augmented Lagrangian approach.
In many real world applications we consider the fair allocation of resources, for example, in course assignments, ad-auctions, and inventory pricing. In [3], the authors consider the auctioning of indivisible goods to multiple strategic buyers in a single-parameter environment. They study EF1, MMS, and NSW fairness conditions. Let there be $n$ buyers and a single seller having $m$ goods to sell. Let $v_{i}\in R_{+}$ be the valuation for buyer $i$ (private value) and let the known public value summarization function among all buyers be $W:2^{[m]}\rightarrow R_{+}$. For a given subset $S\subseteq[m]$, the valuation of the $i^{th}$ buyer is $v_{i}(S)=v_{i}W(S)$. Then, the fairness conditions are defined below.
@@
Motivated by the fairness metrics of [3], [1] proposes an NN model for an auction that has a fair allocation in addition to being IC and IR while maximizing social welfare. The proposed model is described below.